# A non-degeneracy theorem for interacting electrons in one dimension

### v1

@@ 5.1. Perron-Frobenius theorem
(Positivity and reality preserving) Let $L\subset\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$ be a closed subspace and $\mathrm{H}^{1}_{L}(\Omega)$ the space introduced in Definition 3.8. Then we say that
@@
    $\mathrm{H}^{1}_{L}(\Omega)$ is reality preserving if $\mathrm{Re}\,\Psi\in\mathrm{H}^{1}_{L}(\Omega)$ for any $\Psi\in\mathrm{H}^{1}_{L}(\Omega)$.
@@
    $\mathrm{H}^{1}_{L}(\Omega)$ is positivity preserving if $\Psi_{+}=\max\{\Psi,0\}\in\mathrm{H}^{1}_{L}(\Omega)$ for any real-valued $\Psi\in\mathrm{H}^{1}_{L}(\Omega)$.
@@
Let $\Omega\subset\mathbb{R}^{d}$ be open, connected, and bounded subset with Lipschitz boundary, and let $\mathrm{H}^{1}_{L}(\Omega)$ be positivity and reality preserving. Then the Laplacian $-\Delta_{L}$, defined as the unique self-adjoint operator associated to the quadratic form
$a_{0}:\mathrm{H}^{1}_{L}(\Omega)\times\mathrm{H}^{1}_{L}(\Omega)\rightarrow\mathbb{C},\quad a_{0}(\Psi,\Phi)\coloneqq\int_{\Omega}\overline{\nabla\Psi(x)}\bm{\cdot}\nabla\Phi(x)\mathrm{d}x,$
@@
The proof is based on the variational principle and the strong maximum principle. So first, we note that $-\Delta_{L}$ has discrete spectrum and therefore a ground-state $\Psi$ exists. Next, we claim that the ground-state can be taken real-valued. To see this, note that $\overline{\Phi}\in\mathrm{H}^{1}_{L}(\Omega)$ for any $\Phi\in\mathrm{H}^{1}_{L}(\Omega)$ by the reality preserving property. In particular
$a_{0}(\overline{\Psi},\Phi)=\int_{\Omega}\nabla\Psi(x)\bm{\cdot}\nabla\Phi(x)\mathrm{d}x=a_{0}(\overline{\Phi},\Psi)=\lambda\langle\overline{\Phi},\Psi\rangle=\lambda\langle\overline{\Psi},\Phi\rangle,\quad\mbox{for any$\Phi\in\mathrm{H}^{1}_{L}(\Omega)$,}$
where $\lambda\geq 0$ is the ground-state energy. Consequently, $\overline{\Psi}$ is also a ground-state of $-\Delta_{L}$, and therefore $2\mathrm{Re}\Psi=\Psi+\overline{\Psi}$ is a ground-state as well.
Thus, assuming that the ground-state is real-valued, we can define $\Psi_{+}=\max\{0,\Psi\}$ and $\Psi_{-}=(-\Psi)_{+}$. Since $\mathrm{H}^{1}_{L}(\Omega)$ is positivity preserving by assumption, both $\Psi_{+}$ and $\Psi_{-}$ belong to $\mathrm{H}^{1}_{L}(\Omega)$. Moreover, since they have disjoint support and $\Psi=\Psi_{+}-\Psi_{-}$, we have
$a_{0}(\Psi_{+},\Psi_{+})=\int_{\Omega}\overline{\nabla\Psi_{+}(x)}\bm{\cdot}\nabla\Psi_{+}(x)\mathrm{d}x=\int_{\Omega}\overline{\nabla\Psi(x)}\bm{\cdot}\nabla\Psi_{+}(x)\mathrm{d}x=\lambda\langle\Psi,\Psi_{+}\rangle=\lambda\lVert\Psi_{+}\rVert_{\mathrm{L}^{2}}^{2}.$
@@
$\lambda=\min_{\Psi\in\mathrm{H}^{1}_{L}\setminus\{0\}}\frac{a_{0}(\Psi,\Psi)}{\lVert\Psi\rVert_{\mathrm{L}^{2}}^{2}}=\frac{a_{0}(\Psi_{+},\Psi_{+})}{\lVert\Psi_{+}\rVert_{\mathrm{L}^{2}}^{2}}.$
@@
$a_{0}(\Phi,\Psi_{+})=\int_{\Omega}\overline{\nabla\Phi(x)}\bm{\cdot}\nabla\Psi(x)=\lambda\langle\Phi,\Psi_{+}\rangle,\quad\mbox{for any $\Phi\in C_{c}^{\infty}(\Omega)$,}$
@@
Therefore, we can apply the strong maximum principle [Eva10, Theorem 3 in Section 6.4.2] to conclude that either $\Psi_{+}(x)>0$ for any $x\in\Omega$ or $\Psi_{+}=0$. Hence, either $\Psi_{+}=0$ or $\Psi_{-}=0$, which implies $\Psi(x)$ has constant sign everywhere in $\Omega$. Thus any real-valued ground-state of $-\Delta_{L}$ is strictly positive (up to multiplication by a constant). Therefore, the ground-state is non-degenerate as there cannot be two strictly positive functions that are mutually orthogonal. ∎
We can now combine the previous result with the perturbative approach in Reed and Simon [RS78, Section XIII.12] to prove the following Perron-Frobenius theorem for generalized Schrödinger operators. This result will play a key role throughout this paper.
@@
Let $\Omega\subset\mathbb{R}^{d}$ and $\mathrm{H}^{1}_{L}(\Omega)$ be as in Lemma 5.2. Let $V:\mathrm{H}^{1}_{L}(\Omega)\times\mathrm{H}^{1}_{L}(\Omega)\rightarrow\mathbb{C}$ be a sesquilinear form and suppose that there exists a sequence $\{V_{n}\}_{n\in\mathbb{N}}\subset\mathrm{L}^{\infty}(\Omega)$ of real-valued functions such that
$\lim_{n\rightarrow\infty}\lVert V-V_{n}\rVert_{\mathrm{H}^{1}\rightarrow\mathrm{H}^{-1}}=\lim_{n\rightarrow\infty}\sup_{\Psi,\Phi\in\mathrm{H}^{1}_{L}(\Omega)\setminus\{0\}}\frac{\left|V(\Psi,\Phi)-\int_{\Omega}V_{n}(x)\overline{\Psi(x)}\Phi(x)\mathrm{d}x\right|}{\lVert\Psi\rVert_{\mathrm{H}^{1}}\lVert\Phi\rVert_{\mathrm{H}^{1}}}=0.$
Then there exists a unique self-adjoint operator $H(V)=-\Delta_{L}+V$ with quadratic form
$a_{V}(\Psi,\Phi)=\int_{\Omega}\overline{\nabla\Psi(x)}\bm{\cdot}\nabla\Phi(x)\mathrm{d}x+V(\Psi,\Phi),\quad\mbox{for $\Phi,\Psi\in\mathrm{H}^{1}_{L}(\Omega)$.}$ (5.1)
@@
We first show that $V$ is symmetric and $\Delta$-bounded with relative bound 00. For this, note that, for any $\epsilon>0$ there exists $n\in\mathbb{N}$ such that $\lVert V-V_{n}\rVert_{\mathrm{H}^{1}\rightarrow\mathrm{H}^{-1}}\leq\epsilon$. Hence,
$\left|V(\Psi,\Phi)-\int_{\Omega}V_{n}(x)\overline{\Psi(x)}\Phi(x)\mathrm{d}x\right|\leq\epsilon\lVert\Psi\rVert_{\mathrm{H}^{1}}\lVert\Psi\rVert_{\mathrm{H}^{1}}+\lVert V_{n}\rVert_{\mathrm{L}^{\infty}}\lVert\Psi\rVert_{\mathrm{L}^{2}}\lVert\Phi\rVert_{\mathrm{L}^{2}},\quad\mbox{for any $\Psi,\Phi\in\mathrm{H}^{1}_{L}(\Omega)$,}$
@@
As $\epsilon>0$ can be taken arbitrarily small, we conclude that $V$ is symmetric and $\Delta$-bounded with relative bound 00. Hence, by the KLMN theorem (cf. Theorem A.5), there exists a unique self-adjoint operator with quadratic form (5.1). Moreover, as the quadratic form domain of $H(V)$ is $\mathrm{H}^{1}_{L}(\Omega)$, which is compactly embedded in $\mathrm{L}^{2}(\Omega)$, the operator $H(V)$ has compact resolvent and therefore purely discrete spectrum. In particular, there exists at least one ground-state.
To show that the ground-state is non-degenerate and strictly positive almost everywhere, we shall use the following lemma, which is a reformulation of results extracted from Reed and Simon [RS78, Theorem XIII.43 and Theorem XIII.45]. For the proof of this lemma, we refer to their book.
Lemma 5.4 (Perron-Frobenius invariance under strong resolvent convergence via multiplicative potentials).
Let $H$ and $H_{0}$ be self-adjoint operators on $\mathrm{L}^{2}(\Omega)$ and suppose that there exists $\{V_{n}\}_{n\in\mathbb{N}}\subset\mathrm{L}^{\infty}(\Omega)$ such that $H_{0}+V_{n}$ converges in the strong resolvent sense to $H$. Suppose also that $\mathrm{e}^{-tH_{0}}$ is positivity preserving (i.e., sends non-negative functions to non-negative functions). Then the following holds: if $H_{0}$ has a non-degenerate and almost everywhere strictly positive ground-state, then so does $H$ (provided that a ground-state exists).
We now observe that, by Lemma 5.2, the operator $H_{0}=-\Delta_{L}$ has a non-degenerate and strictly positive ground-state. Moreover, $\mathrm{e}^{-tH_{0}}$ is positivity preserving because, from the standard formula
$\nabla|\Psi|=\nabla\Psi\mathbb{1}_{\{\Psi>0\}}-\nabla\Psi\mathbb{1}_{\{\Psi<0\}},\quad\mbox{where $\mathbb{1}_{A}$ is the indicator function on$A$,}$
for real-valued $\Psi$, we see that $H_{0}$ satisfies the Beurling-Deny criterion (cf. [RS78, Theorem XIII.50]):
$\langle|\Psi|,H_{0}|\Psi|\rangle=\int_{\Omega}|\nabla|\Psi|(x)|^{2}\mathrm{d}x=\int_{\Omega}|\nabla\Psi(x)|^{2}\mathrm{d}x=\langle\Psi,H\Psi\rangle,\quad\mbox{for any $\Psi\in\mathrm{H}^{1}_{L}(\Omega)$.}$
To conclude, we now note that, since $V_{n}\rightarrow V$ in $\mathcal{B}(\mathrm{H}^{1};\mathrm{H}^{-1})$, the operators $H_{0}+V_{n}$ converge to $H(V)$ in the strong (in fact norm) resolvent sense. Indeed, this follows from the resolvent formula
@@ 5.2. Unitary reduction to the simplex
The key observation that allow us to apply Theorem 5.3 to the operator $H_{N}^{L}(v,w)$ is that the hypercube $I_{N}$ can be decomposed in a disjoint union of reflections of the simplex $S_{N}$. To make this precise, let us introduce some additional notation.
Figure 1. Simplex $S_{N}$ in the case $N=2$ (left) and $N=3$ (right) with edges of $\partial I_{N}$ (dashed lines) and interior boundary $\Gamma_{\rm int}\subset\partial S_{N}$ (in red).
@@
$\mathrm{H}^{1}_{L}(S_{N})\coloneqq\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}_{L}(I_{N})\cap\mathcal{H}_{N}\}.$ (5.5)
We then have the following key decomposition of $\mathcal{H}_{N}\cap\mathrm{H}^{1}_{L}(I_{N})$.
@@
The map $T:\mathrm{H}^{1}_{L}(S_{N})\rightarrow\mathrm{H}^{1}_{L}(I_{N})\cap\mathcal{H}_{N}$ defined via
@@
The fact that $\sigma(\overline{S_{N}})$ covers $I_{N}$ is clear. Indeed, for any $(x_{1},...,x_{N})\in I_{N}$ there exists $\sigma\in\mathcal{P}_{N}$ such that $0\leq x_{\sigma(1)}\leq x_{\sigma(2)}...\leq x_{\sigma(N)}\leq 1$ and therefore $(x_{1},...,x_{N})\in\sigma^{-1}(\overline{S_{N}})$.
@@
We now turn to (ii). First, it is clear from the definition (5.3) that $F_{0}\cup F_{N}\subset\partial I_{N}$ and therefore $(F_{0}\cup F_{N})\cap I_{N}=\emptyset$. Hence, $\partial S_{N}\cap I_{N}\subset\Gamma_{\rm int}$. On the other hand, if $(x_{1},..,x_{N})\in\Gamma_{\rm int}=\partial S_{N}\setminus(F_{0}\cup F_{N})$ we must have $10<x_{0}\leq x_{1}...\leq x_{N}<$ 1. As every element in the boundary of $I_{N}$ must have at least one coordinate with value 00 or $1$, we must have $(x_{1},...,x_{N})\in I_{N}$ and therefore item (ii) holds.
We now show that $T$ is an isometry from $\mathrm{H}^{1}_{L}(S_{N})$ to $\mathrm{H}^{1}(I_{N})\cap\mathcal{H}_{N}$. To this end, first note that, since $\Psi$ is the restriction of an antisymmetric function, it follows from the definition of $\Gamma_{\rm int}$ (see (5.4)) that $\gamma\Psi\rvert_{\Gamma_{\rm int}}=0$. Consequently, from property (ii), the extension of $\Psi$ to $I_{N}$ by zero on $I_{N}\setminus S_{N}$ belongs to $\mathrm{H}^{1}(I_{N})$. Hence $T\Psi\in\mathrm{H}^{1}(I_{N})$. Moreover, if we denote the permutation of $x_{j}$ and $x_{j+1}$ by $\sigma_{j}$, then by construction we have
$(\sigma_{j})_{\#}(T\Psi)=\frac{1}{\sqrt{N!}}\sum_{\sigma\in\mathcal{P}_{N}}\operatorname{sign}(\sigma)(\sigma_{j}\circ\sigma)_{\#}(\Psi)=-\frac{1}{\sqrt{N!}}\sum_{\sigma_{j}\circ\sigma\in\mathcal{P}_{N}}\mathrm{sgn}(\sigma_{j}\circ\sigma)(\sigma_{j}\circ\sigma)_{\#}(\Psi)=-T\Psi.$

### v2

@@ 4.1. Perron-Frobenius theorem
(Positivity and reality preserving) Let $B\subset\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$ be a closed subspace and $\mathrm{H}^{1}_{B}(\Omega)$ the space introduced in Definition 3.8. Then we say that
@@
    $\mathrm{H}^{1}_{B}(\Omega)$ is reality preserving if $\mathrm{Re}\,\Psi\in\mathrm{H}^{1}_{B}(\Omega)$ for any $\Psi\in\mathrm{H}^{1}_{B}(\Omega)$.
@@
    $\mathrm{H}^{1}_{B}(\Omega)$ is positivity preserving if $\Psi_{+}=\max\{\Psi,0\}\in\mathrm{H}^{1}_{B}(\Omega)$ for any real-valued $\Psi\in\mathrm{H}^{1}_{B}(\Omega)$.
@@
Let $\Omega\subset\mathbb{R}^{d}$ be open, connected, and bounded subset with Lipschitz boundary, and let $\mathrm{H}^{1}_{B}(\Omega)$ be positivity and reality preserving. Then the Laplacian $\Delta_{B}$, defined as the unique self-adjoint operator associated to the quadratic form
$\mathfrak{a}_{0}:\mathrm{H}^{1}_{B}(\Omega)\times\mathrm{H}^{1}_{B}(\Omega)\rightarrow\mathbb{C},\quad\mathfrak{a}_{0}(\Psi,\Phi)\coloneqq\int_{\Omega}\overline{\nabla\Psi(x)}\boldsymbol{\cdot}\nabla\Phi(x)\mathrm{d}x,$
@@
Remark.
The above result appears in various forms throughout the literature [Eva10, Chapter 6], and the general form stated above is likely known to experts. Nevertheless, due to the lack of a reference for this precise statement, we provide the simple proof below.
@@
The proof is based on the variational principle and the strong maximum principle. So first, we note that $\Delta_{B}$ has discrete spectrum and therefore a ground-state $\Psi$ exists. Next, we claim that the ground-state can be taken real-valued. To see this, note that $\overline{\Phi}\in\mathrm{H}^{1}_{B}(\Omega)$ for any $\Phi\in\mathrm{H}^{1}_{B}(\Omega)$ by the reality preserving property. In particular
$\mathfrak{a}_{0}(\overline{\Psi},\Phi)=\int_{\Omega}\nabla\Psi(x)\boldsymbol{\cdot}\nabla\Phi(x)\mathrm{d}x=\mathfrak{a}_{0}(\overline{\Phi},\Psi)=\lambda\langle\overline{\Phi},\Psi\rangle=\lambda\langle\overline{\Psi},\Phi\rangle,\quad\mbox{for any $\Phi\in\mathrm{H}^{1}_{B}(\Omega)$,}$
where $\lambda\geq 0$ is the ground-state energy. Consequently, $\overline{\Psi}$ is also a ground-state of $\Delta_{B}$, and therefore $2\mathrm{Re}\Psi=\Psi+\overline{\Psi}$ is a ground-state as well.
Thus, assuming that the ground-state is real-valued, we can define $\Psi_{+}=\max\{0,\Psi\}$ and $\Psi_{-}=(-\Psi)_{+}$. Since $\mathrm{H}^{1}_{B}(\Omega)$ is positivity preserving by assumption, both $\Psi_{+}$ and $\Psi_{-}$ belong to $\mathrm{H}^{1}_{B}(\Omega)$. Moreover, since they have disjoint support and $\Psi=\Psi_{+}-\Psi_{-}$, we have
$\mathfrak{a}_{0}(\Psi_{+},\Psi_{+})=\int_{\Omega}\overline{\nabla\Psi_{+}(x)}\boldsymbol{\cdot}\nabla\Psi_{+}(x)\mathrm{d}x=\int_{\Omega}\overline{\nabla\Psi(x)}\boldsymbol{\cdot}\nabla\Psi_{+}(x)\mathrm{d}x=\lambda\langle\Psi,\Psi_{+}\rangle=\lambda\lVert\Psi_{+}\rVert_{\mathrm{L}^{2}}^{2}.$
@@
$\lambda=\min_{\Psi\in\mathrm{H}^{1}_{B}\setminus\{0\}}\frac{\mathfrak{a}_{0}(\Psi,\Psi)}{\lVert\Psi\rVert_{\mathrm{L}^{2}}^{2}}=\frac{\mathfrak{a}_{0}(\Psi_{+},\Psi_{+})}{\lVert\Psi_{+}\rVert_{\mathrm{L}^{2}}^{2}}.$
@@
$\mathfrak{a}_{0}(\Phi,\Psi_{+})=\int_{\Omega}\overline{\nabla\Phi(x)}\boldsymbol{\cdot}\nabla\Psi(x)=\lambda\langle\Phi,\Psi_{+}\rangle,\quad\mbox{for any $\Phi\in C_{c}^{\infty}(\Omega)$,}$
@@
Therefore, we can apply the strong maximum principle [Eva10, Theorem 3 in Section 6.4.2] to conclude that either $\Psi_{+}(x)>0$ for any $x\in\Omega$ or $\Psi_{+}=0$. Hence, either $\Psi_{+}=0$ or $\Psi_{-}=0$, which implies $\Psi(x)$ has constant sign everywhere in $\Omega$. Thus any real-valued ground-state of $\Delta_{B}$ is strictly positive (up to multiplication by a constant). As a consequence, the ground-state is non-degenerate as there cannot be two strictly positive functions that are mutually orthogonal. ∎
Our next goal is to combine the previous result with the perturbative approach in Reed and Simon [RS78, Section XIII.12] to prove a Perron-Frobenius theorem for Schrödinger operators with generalized potentials. More precisely, we shall use the following lemma, which is a reformulation of results extracted from Reed and Simon [RS78, Theorem XIII.43 and Theorem XIII.45]. For the proof of this lemma, we refer to their book.
Lemma 4.3 (Non-degeneracy via perturbation).
Let $H$ and $H_{0}$ be self-adjoint operators on $\mathrm{L}^{2}(\Omega)$. Suppose the following holds:
  1. (i)
    (Positive semigroup) The heat semigroup $\mathrm{e}^{-tH_{0}}$ is positivity preserving (i.e., sends non-negative functions to non-negative functions).
  2. (ii)
    (Unperturbed positive ground-state) $H_{0}$ has a non-degenerate and almost everywhere strictly positive ground-state.
  3. (iii)
    (Approximation property) There exists $\{V_{n}\}_{n\in\mathbb{N}}\subset\mathrm{L}^{\infty}(\Omega)$ such that $H_{0}+V_{n}$ converges in the strong resolvent sense to $H$.
Then $H$ has a non-degenerate and almost everywhere strictly positive ground-state (provided that a ground-state exists).
Next, let us recall the Beurling-Deny criterion to establish positivity of the heat semigroup [RS78, Theorem XIII.50]. Again, we refer to [RS78] for the proof.
Lemma 4.4 (Beurling-Deny criterion).
Let $H\geq 0$ be a self-adjoint operator on $\mathrm{L}^{2}(\Omega)$ and $\mathfrak{a}_{H}:\mathcal{Q}(H)\rightarrow\mathbb{R}$ be the associated quadratic form. Then the following are equivalent:
  1. (a)
    The heat semigroup $e^{-tH}$ is positivty preserving, i.e., sends non-negative functions to non-negative functions.
  2. (b)
    $e^{-tH}$ is reality preserving, the form domain $\mathcal{Q}(H)$ is positivity preserving, and
    $\mathfrak{a}_{H}(u_{+})+\mathfrak{a}_{H}(u_{-})\leq\mathfrak{a}_{H}(u),\quad\mbox{for any real-valued $u\in\mathcal{Q}(H)$.}$
  3. (c)
    $\mathcal{Q}(H)$ is absolute value preserving, i.e., $|\Psi|\in\mathcal{Q}(H)$ for any $\Psi\in\mathcal{Q}(H)$, and
    $\mathfrak{a}_{H}(|\Psi|)\leq\mathfrak{a}_{H}(\Psi),\quad\mbox{for any $\Psi\in\mathcal{Q}(H)$.}$
Remark 4.5 (Reality and posivity preserving vs. absolute value preseving).
It is interesting to note that any closed subspace $X\subset\mathrm{H}^{1}(\Omega)$ is positivity and reality preserving if and only if it is absolute value preserving. To be precise, the if direction follows by noticing that
$|\alpha\Psi+|\Psi||=\sqrt{(\alpha\mathrm{Re}\,\Psi+|\Psi|)^{2}+\alpha^{2}(\mathrm{Im}\,\Psi)^{2}}=\sqrt{(1+\alpha^{2})|\Psi|^{2}+2\alpha\mathrm{Re}\,\Psi|\Psi|}\in X$
and taking the limit $\lim_{\alpha\downarrow 0}\frac{|\alpha\Psi+|\Psi||-|\Psi|}{\alpha}=\mathrm{Re}\Psi$ to conclude that $X$ is reality preserving. The other direction follows by defining the Laplacian with form domain $X$ and using $(b)\Rightarrow(c)$ in Lemma 4.4.
We can now combine the previous lemmas to prove the following version of the Perron-Frobenius theorem.
@@
Let $\Omega\subset\mathbb{R}^{m}$ and $\mathrm{H}^{1}_{B}(\Omega)$ be as in Lemma 4.2. Let $V:\mathrm{H}^{1}_{B}(\Omega)\times\mathrm{H}^{1}_{B}(\Omega)\rightarrow\mathbb{C}$ be a sesquilinear form and suppose that there exists a sequence $\{V_{n}\}_{n\in\mathbb{N}}\subset\mathrm{L}^{\infty}(\Omega)$ of real-valued functions such that
$\lim_{n\rightarrow\infty}\lVert V-V_{n}\rVert_{\mathrm{H}^{1}\rightarrow\mathrm{H}^{-1}}=\lim_{n\rightarrow\infty}\sup_{\Psi,\Phi\in\mathrm{H}^{1}_{B}(\Omega)\setminus\{0\}}\frac{\left|V(\Psi,\Phi)-\int_{\Omega}V_{n}(x)\overline{\Psi(x)}\Phi(x)\mathrm{d}x\right|}{\lVert\Psi\rVert_{\mathrm{H}^{1}}\lVert\Phi\rVert_{\mathrm{H}^{1}}}=0.$
Then there exists a unique self-adjoint operator $H(V)=-\Delta_{B}+V$ with quadratic form
$\mathfrak{a}_{V}(\Psi,\Phi)=\int_{\Omega}\overline{\nabla\Psi(x)}\boldsymbol{\cdot}\nabla\Phi(x)\mathrm{d}x+V(\Psi,\Phi),\quad\mbox{for $\Phi,\Psi\in\mathrm{H}^{1}_{B}(\Omega)$.}$ (4.1)
@@
We first show that $V$ is symmetric and $\Delta$-bounded with relative bound 0. For this, note that, for any $\epsilon>0$ there exists $n\in\mathbb{N}$ such that $\lVert V-V_{n}\rVert_{\mathrm{H}^{1}\rightarrow\mathrm{H}^{-1}}\leq\epsilon$. Hence,
$\left|V(\Psi,\Phi)\right|\leq\epsilon\lVert\Psi\rVert_{\mathrm{H}^{1}}\lVert\Phi\rVert_{\mathrm{H}^{1}}+\lVert V_{n}\rVert_{\mathrm{L}^{\infty}}\lVert\Psi\rVert_{\mathrm{L}^{2}}\lVert\Phi\rVert_{\mathrm{L}^{2}},\quad\mbox{for any $\Psi,\Phi\in\mathrm{H}^{1}_{B}(\Omega)$,}$
@@
As $\epsilon>0$ can be taken arbitrarily small, we conclude that $V$ is symmetric and $\Delta$-bounded with relative bound 0. Hence, by the KLMN theorem (cf. [RS75, Theorem X.17]), there exists a unique self-adjoint operator with quadratic form (4.1). Moreover, as the quadratic form domain of $H(V)$ is $\mathrm{H}^{1}_{B}(\Omega)$, which is compactly embedded in $\mathrm{L}^{2}(\Omega)$, the operator $H(V)$ has compact resolvent and therefore purely discrete spectrum. In particular, there exists at least one ground-state.
To show that the ground-state is non-degenerate and strictly positive almost everywhere, we can now use Lemma 4.3. For this, we set $H_{0}\coloneqq-\Delta_{B}$, and note that $H_{0}$ has a non-degenerate and strictly positive ground-state by Lemma 4.2. Moreover, since $H_{0}$ is a real operator, i.e., commutes with complex conjugation, the semigroup $e^{-tH_{0}}$ is reality preserving. Hence, one can easily verify that criterion (b) in Lemma 4.4 holds for $H_{0}$, and therefore, $\mathrm{e}^{-tH_{0}}$ is also positivity preserving. In particular, assumptions (i) and (ii) from Lemma 4.3 hold. To verify assumption (iii), note that, since $V_{n}\rightarrow V$ in $\mathcal{B}(\mathrm{H}^{1};\mathrm{H}^{-1})$, the operators $H_{0}+V_{n}$ converge to $H(V)$ in the strong (in fact norm) resolvent sense. Indeed, this follows from the formula
@@ 4.2. Unitary reduction to the simplex
The key observation that allow us to apply Theorem 4.6 to the operator $H_{N}^{B}(v,w)$ is that the hypercube $I_{N}$ can be decomposed in a disjoint union of reflections of the simplex $S_{N}$. To make this precise, let us introduce some additional notation.
Figure 1. Simplex $S_{N}$ in the case $N=2$ (left) and $N=3$ (right) with edges of $\partial I_{N}$ (dashed lines) and interior boundary $\Gamma_{\rm int}\subset\partial S_{N}$ (in red), where functions in $\mathcal{Q}_{N}(B;S_{N})$ vanish.
@@
$\mathcal{Q}_{N}(B;S_{N})\coloneqq\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}_{B}(I_{N})\cap\mathcal{H}_{N}\}.$ (4.5)
Note that any $\Psi\in\mathrm{H}^{1}(I_{N})\cap\mathcal{H}_{N}$ has a well-defined and identically zero (Dirichlet) trace along the coincidence hyperplanes $\{x\in I_{N}:x_{j}=x_{k}\mbox{ for $j\neq k$}\}$ by antisymmetry. In particular, any $\Psi\in\mathcal{Q}_{N}(B;S_{N})$ vanishes along the interior boundary $\Gamma^{\rm int}$.
Using this observation, we then have the following key decomposition of $\mathcal{Q}_{N}(B)=\mathcal{H}_{N}\cap\mathrm{H}^{1}_{B}(I_{N})$.
@@
The map $T:\mathcal{Q}_{N}(B;S_{N})\rightarrow\mathcal{Q}_{N}(B)$ defined via
@@
The fact that $\cup_{\sigma}\sigma(\overline{S_{N}})$ covers $I_{N}$ is clear. Indeed, for any $(x_{1},...,x_{N})\in I_{N}$ there exists $\sigma\in\mathcal{P}_{N}$ such that $0\leq x_{\sigma(1)}\leq x_{\sigma(2)}...\leq x_{\sigma(N)}\leq 1$ and therefore $(x_{1},...,x_{N})\in\sigma^{-1}(\overline{S_{N}})$.
@@
We now turn to (ii). First, it is clear from the definition (4.3) that $F_{0}\cup F_{N}\subset\partial I_{N}$ and therefore $(F_{0}\cup F_{N})\cap I_{N}=\emptyset$. Hence, $\partial S_{N}\cap I_{N}\subset\Gamma_{\rm int}$. On the other hand, if $(x_{1},..,x_{N})\in\Gamma_{\rm int}=\partial S_{N}\setminus(F_{0}\cup F_{N})$ we must have $0<x_{0}\leq x_{1}...\leq x_{N}<1$. As every element in the boundary of $I_{N}$ must have at least one coordinate with value 0 or $1$, we must have $(x_{1},...,x_{N})\in I_{N}$ and therefore item (ii) holds.
We now show that $T$ is an isometry from $\mathcal{Q}_{N}(B;S_{N})$ to $\mathcal{Q}_{N}(B)$. To this end, first note that, since $\Psi$ is the restriction of an antisymmetric function, it follows from the definition of $\Gamma_{\rm int}$ (see (4.4)) that $\gamma\Psi\rvert_{\Gamma_{\rm int}}=0$. Consequently, from property (ii), the extension of $\Psi$ to $I_{N}$ by zero on $I_{N}\setminus S_{N}$ belongs to $\mathrm{H}^{1}(I_{N})$. Hence $T\Psi\in\mathrm{H}^{1}(I_{N})$. Moreover, if we denote the permutation of $x_{j}$ and $x_{j+1}$ by $\sigma_{j}$, then by construction we have
$(\sigma_{j})_{\#}(T\Psi)=\frac{1}{\sqrt{N!}}\sum_{\sigma\in\mathcal{P}_{N}}\operatorname{sgn}(\sigma)(\sigma_{j}\circ\sigma)_{\#}(\Psi)=-\frac{1}{\sqrt{N!}}\sum_{\sigma_{j}\circ\sigma\in\mathcal{P}_{N}}\mathrm{sgn}(\sigma_{j}\circ\sigma)(\sigma_{j}\circ\sigma)_{\#}(\Psi)=-T\Psi.$